# ExploraCoder: Advancing code generation for multiple unseen APIs via planning and chained exploration

### v1

@@ 4 Benchmark Construction
Unseen library benchmarks provide essential references for evaluating LLMs’ performance in handling unseen APIs. Real-world programming often involves multiple unseen APIs. However, existing benchmarks typically provide problems with only 1-2 APIs. To address this limitation, we constructed two Torchdata-based [^3] Torchdata is released in May 2022. The LLMs used in this study (GPT-3.5-turbo and GPT-4-0613) were pretrained on GitHub corpora before Torchdata’s release, meaning the APIs were unseen by them during training. Moreover, since its release, Torchdata has accumulated open-source client code, allowing more recently trained LLMs to acquire knowledge of Torchdata. The choice of Torchdata library provides an opportunity to compare our training-free framework with models containing pretrained knowledge. multi-API benchmarks: Torchdata-Github and Torchdata-Manual. Detailed statistics are presented in Table 1.
Table 1: Statistical Summary of two Torchdata-based benchmarks. Num. APIs. and Avg. APIs. report the range and average number of distinct APIs involved in each sample. Num. Invoc. and Avg. Invoc. report the range and average number of API invocations in the samples’ canonical solution. Volume of the doc pool refers to the number of API documents provided by the library, which also represents the size of the search space during API retrieval.
Benchmarks
Num.
samples
Num.
APIs
Avg.
APIs
Num.
Invoc.
Avg.
Invoc.
Volume of
doc pool
Torchdata-Github 50 3-8 4.26 3-8 4.64 228
Torchdata-Manual 50 8-14 9.94 8-21 12.00 228
Torchdata-Github. We curated the 50 Torchdata problems provided by Ma et al. (2024b) and constructed the Torchdata-Github benchmark. These programming problems are adapted from client project of Torchdata on GitHub, each containing a coarse-grained user requirement that entails 3-8 API invocations. Specifically, we add example inputs for each problem to facilitate CoAE. And we manualy supplemented external resources needed to run test cases in some problems [^4] Some external resources, such as local files to be loaded in problems, are not provided by Ma et al. (2024b).
Torchdata-Manual. To evaluate the models’ ability to address more complex API invocations, we developed a new benchmark called Torchdata-Manual, comprising 50 manually crafted programming problems. Each problem involves 8-14 distinct Torchdata APIs and more clearly stated problem descriptions. To ensure the diversity of the programming tasks, we randomly sampled numerous API combinations from the Torchdata documentation pool and selected plausible combinations to formulate the problem. Two programmers with more than five years of Python coding experience are invited to review the benchmark. More detailed construction methodology is provided in the Appendix A.8.
@@ 5.1 Experimental setups
Benchmarks and base language models. We evaluate ExploraCoder’s performance on multi-API invocation problems using the Torchdata-Github and Torchdata-Manual benchmarks. Based on the the publicly available information on models’ training data cutoff date, we conduct our main experiments under two base models settings: (1) API-untrained model: where the API knowledge is unseen by model during training phase. We choose GPT-3.5-turbo-0125 and GPT-4-0613555See gpt-3.5-turbo and gpt-4 on https://platform.openai.com/docs/models/overview as representatives. (2) API-pretrained model: where the API knowledge is pretrained in model. We represent it by a newer version of GPT-4-1106-preview and two SOTA opensource code LLM: CodeQwen-1.5 and DeepseekCoder-6.7b [^6] We use the instruct-tuned version of code LLM. See https://qwenlm.github.io/blog/codeqwen1.5/ and https://deepseekcoder.github.io/. Due to the token budgets and inference speed, we primarily experiment ExploraCoder with GPT-3.5-turbo-0125, while reporting GPT-4-0613 results where necessary to further support our conclusions.
Evaluation metrics. We adopt Pass@k as our primary evaluation metrics in accordance with previous work (Zan et al., 2023). For each problem, we randomly sample $n\geq k$ code solutions from the model to execute against test cases. And pass@k is calculated as the percentage of problems solved using k candidates per problem.
In our preliminary study with Torchdata-Manual, we find that some baselines cannot pass any programming problem at all. To better observe their nuance performance differences, we additionally report Success@k (Chen et al., 2024) in experiments. Success rate relaxes the evaluation criteria by measuring whether the generated code can be executed successfully without runtime errors within limited timeout constraints.
Implementation details. We implement ExplorCoder by setting $k_{\mathcal{D}}=5$ for task planning module. For API recommendation, we set $k=20$ as initial retrieval volume, and we set $k_{G}=15$ on Torchdata-Github following Ma et al. (2024b) and $k_{G}=20$ on Torchdata-Manual. For CoAE, we set $m=5$. To generate diverse candidates, we set the $temperature=0.8$ and $top\_p=0.95$ for our CoAE and for final solution generation across all baselines. More detailed experimental settings are left in Appendix A.9
5.2 Generating multiple API invocations with LLMs
Firstly, we analyze the effectiveness of different knowledge injection approaches in library-oriented code generation tasks in Table 2 and 3.
Invoking APIs using API-untrained and API-pretrained models. By analyzing the direct generation performance of the five base models, we observe that API-pretrained models consistently outperform API-untrained models. This highlights the importance of prior API knowledge in library-oriented code generation. And the lower performance across all models on the Torchdata-Manual (Table 3) further underscores the challenge posed by more complex API invocations, making it a more effective benchmark for our task.  [^7] Notably, API-untrained models manage to solve a small number of problems in Torchdata-Github even without explicit API information in the context. We find in Appendix A.10 that these models circumvent the lack of API knowledge by invoking built-in APIs that provide similar functionalities and meet the task requirements.
Table 2: Performance comparison on Torchdata-Github. We compare with SOTA approach in later section.
API Knowledge Method $k=1$ $k=5$ $k=10$ $k=20$
Pass Success Pass Success Pass Success Pass Success
@@ 5.2 Generating multiple API invocations with LLMs
Table 3: Performance comparison on Torchdata-Manual. We compare with SOTA approach in later section
API Knowledge Method $k=1$ $k=5$ $k=10$ $k=20$
Pass Success Pass Success Pass Success Pass Success
@@
GPT-4-1106-preview 0.19% 1.33% 0.95% 6.15% 1.90% 11.41% 3.80% 21.05%
+ naive RAG 1.43% 3.90% 5.87% 15.32% 9.51% 23.03% 13.90% 27.90%
+ ExploraCoder 9.91% 29.86% 23.66% 51.76% 29.30% 58.19% 33.71% 63.10%
@@
+ naive RAG 0.10% 0.48% 0.48% 2.20% 0.95% 3.90% 1.90% 5.90%
+ ExploraCoder 4.48% 9.71% 8.71% 18.60% 11.61% 22.45% 13.79% 25.80%
GPT-4-0613 0% 0% 0% 0% 0% 0% 0% 0%
+ naive RAG 0% 0.57% 0% 2.59% 0% 4.61% 0% 7.71%
+ ExploraCoder 10.85% 21.24% 19.59% 35.73% 23.27% 40.87% 27.71% 45.70%
@@
Through a naive RAG framework (Zhou et al., 2023), the performance of API-untrained models has been effectively improved, bridging the gap caused by the lack of API knowledge compared to API-pretrained models. Specifically, GPT-3.5-turbo-0125 + naive RAG achieves 7.36% absolute increase in pass@20 compare to its base model across two datasets. We make a fairer comparison of RAG methods and pretraining methods by looking into two GPT-4 models. We discussed the fairness in Appendix A.2. GPT-4-0613 + naive RAG outperforms GPT-4-1106-preview by an average of 2.69% absolute increase in pass@k on Torchdata-Github, but underperforms on the more challenging Torchdata-Manual benchmark.
@@
ExploraCoder vs naive RAG on API-untrained models. From Table 2 and 3, we can observe that ExploraCoder brings substantial improvements over naive RAG for both two API-untrained models (GPT-3.5-turbo-0125 and GPT-4-0613), with an average absolute gains in pass@20 of 3.5% on Torchdata-Github and 19.8% on Torchdata-Manual. These improvements could be attributed to ExploraCoder’s potential in addressing two limitations of the naive RAG framework when handling complex API invocation subtasks:
@@
(1) Retrieval for complex requirement: In the naive RAG approach, the retriever’s ability to recall relevant APIs for comprehensive requirements becomes a bottleneck (Please refer to Appendix A.7). ExploraCoder addresses this by using a divide-and-conquer strategy, retrieving APIs for each individual invocation subtask. Additionally, ExploraCoder alleviates dependency on the retrieval number hyperparameter by setting a fixed API number for simple subtasks and dynamically adjusting the subtask number.
(2) Generating code with multiple unseen APIs: Current models struggle to generate code that invokes multiple unseen APIs due to the cognitive complexity involved in understanding new APIs and managing their interactions (Please refer to Appendix A.10 for case study). ExploraCoder mitigates this challenge by adopting a human-like exploratory programming paradigm, where it incrementally generates simple, reusable code snippets that are highly relevant to problem solving during CoAE, and learns the API usage knowledge by observing the API interactions behavior.
ExploraCoder on API-pretrained model. We observe that the API-pretrained GPT-4-1106-preview underperforms on Torchdata-Manual, achieving a pass@1 of only 0.19%. Therefore, we use GPT-4-1106-preview on Torchdata-Manual benchmark as a proxy to further examine the effectiveness of ExploraCoder on API-pretrained models. As shown in Table 3, ExploraCoder achieves a substantial improvement over GPT-4-1106-preview, with an absolute pass@1 increase of 9.72%, and it also outperforms GPT-4-1106-preview + naive RAG by 8.48%. These results indicate that ExploraCoder is universally effective, improving models with varying levels of pretraining on relevant API knowledge.
5.3 Boosting ExploraCoder with experience exploitation
Table 4: Performance comparison between ExploraCoder and ExploraCoder* on Torchdata-Manual.
Benchmark Model Method $k=1$ $k=5$ $k=10$ $k=20$
Pass Success Pass Success Pass Success Pass Success
Torchdata-Github GPT-3.5-turbo ExploraCoder 10.19% 19.50% 18.64% 39.39% 21.67% 48.56% 25.62% 57.30%
ExploraCoder* 19.24% 38.66% 25,41% 54.93% 27.64% 59.56% 31.62% 63.71%
GPT-4-0613 ExploraCoder 15.43% 23.10% 21.53% 45.62% 28.11% 55.25% 30.00% 61.87%
ExploraCoder* 19.52% 40.19% 25.50% 58.78% 27.66% 68.05% 31.18% 70.37%
Torchdata-Manual GPT-3.5-turbo ExploraCoder 4.48% 9.71% 8.71% 18.60% 11.61% 22.45% 13.79% 25.80%
ExploraCoder* 8.76% 15.48% 14.04% 25.31% 16.75% 29.91% 19.81% 33.81%
GPT-4-0613 ExploraCoder 10.85% 21.24% 19.59% 35.73% 23.27% 40.87% 27.71% 45.70%
ExploraCoder* 15.24% 30.00% 29.77% 54.58% 36.22% 62.71% 41.70% 69.52%
Our quantitative analysis (Please refer to Appendix A.5) shows a positive correlation between CoAE success rate and the final solution success rate. In cases that have more successful executed subtasks, particularly those with a CoAE success rate of 1, are more likely to generate successful or passing final solutions. Intuitively, the failures in the exploration chain, such as wrong API usage could propagate through API interactions and affect the generation of final solutions. Therefore, we wonder if we can enhance ExploraCoder by improving the success rate in each API invocation subtask.
To this end, we designed an enhanced ExploraCoder* by introducing an extra self-debug step into the CoAE. When the exploration tries reaches a threshold number and no candidate code is executable, instead of keep exploring more candidates codes, we enhance the exploitation of the existing candidates by prompting LLMs to debug on the failed ones. Table 4 shows that ExploraCoder* significantly boosts the final solution’s quality, achieving an average increase of 50.33% in success@1 and 68% in pass@1.
5.4 Comparing with related approaches
In this section, we further compare ExploraCoder with other advanced RAG-based approaches. We also include Docprompting, the previously reported naive RAG framework, along with direct generation method as baselines.
For CAPIR and EpiGen, we set a fixed number API recommendation in accordance with our $A_{G}$, and we directly use the subtasks generated by ExploraCoder’s planning module as the CoT instruction for EpiGen. Tables 5 and 6 show that both CAPIR and EpiGen improve upon the naive RAG baseline, while ExploraCoder further surpasses these two approaches, with an relative increase of 82.6% on pass@1 across two benchmarks.
To enable a fair comparison with ExploraCoder*, we adapted a SOTA debugging framework, Self-Repair (Olausson et al., 2024), with retrieval augmentation. We incorporate CAPIR’s API retrieval results into the LLM’s context throughout Self-Repair’s 3-stage generation process. We compare ExploraCoder and Self-Repair under the same computational budget. Specifically, for each problem, ExploraCoder generates $n$ plans, enabling up to $n$ debug operations in CoAE, we set the iteration budget for Self-repair to $n$ accordingly.
Table 5: Comparing ExploraCoder with baseline approaches using GPT3.5 on Torchdata-Github.
@@ 5.4 Comparing with related approaches
Direct 1.70% 2.09% 5.54% 6.95% 7.28% 9.64% 8.00% 11.90%
DocPrompting (Zhou et al., 2023) 6.00% 10.57% 10.55% 24.00% 14.67% 32.50% 20.83% 40.81%
CAPIR (Ma et al., 2024b) 5.90% 10.47% 14.59% 27.08% 18.60% 37.19% 23.52% 47.43%
EpiGen (Li et al., 2024) 8.57% 18.95% 14.63% 35.61% 17.24% 41.67% 19.61% 47.62%
Self-Repair(Olausson et al., 2024) 16.47% 22.10% 21.04% 29.70% 21.75% 32.20% 22.00% 33.90%
ExploraCoder* (Ours) 19.24% 38.66% 25,41% 54.93% 27.64% 59.56% 31.62% 63.71%
Table 6: Comparing ExploraCoder with baseline approaches using GPT3.5 on Torchdata-Manual.
@@
Direct 0% 0% 0% 0% 0% 0% 0% 0%
DocPrompting (Zhou et al., 2023) 0.10% 0.48% 0.48% 2.20% 0.95% 3.90% 1.90% 5.90%
CAPIR (Ma et al., 2024b) 2.76% 3.81% 5.17% 9.51% 5.79% 12.77% 6.00% 15.81%
EpiGen (Li et al., 2024) 1.62% 5.73% 3.16% 13.44% 3.75% 17.00% 3.97% 21.43%
ExploraCoder (Ours) 4.48% 9.71% 8.71% 18.60% 11.61% 22.45% 13.79% 25.80%
Self-Repair(Olausson et al., 2024) 5.33% 14.19% 6.84% 18.30% 7.48% 19.66% 7.99% 20.00%
ExploraCoder* (Ours) 8.76% 15.48% 14.04% 25.31% 16.75% 29.91% 19.81% 33.81%
Table 5 and 6 shows that ExploraCoder outperforms self-repair by a significant margin of 40.6% on pass@1. While Self-Repair can effectively boost the success rate for CAPIR, the overall improvement in the pass rate remains limited. This limitation stems from the nature of next-token prediction models, where API usage hallucinations tend to accumulate throughout the API invocation sequence, particularly in subtasks involving multiple APIs. In many cases, we observe that Self-Repair repeatedly attempts to fix a buggy codes that deviates substantially from the correct solution, continuing until it exhausts its repair budget (Please refer to Appendix A.10 for case study). In contrast, ExploraCoder mitigates this issue by leveraging timely execution feedback during simpler early-stage tasks, using this feedback to inform and generate accurate API invocations. Additionally, Self-Repair demonstrates significant performance gains when considering a small number of output candidates. However, its relative improvement decreases as k increases. This is because the debugging process in Self-Repair relies on explicit and consistent feedback, leading to highly similar solutions. ExploraCoder can better harness the potential of output candidates by utilizing a standalone solution generator.

### v2

@@ 4 Benchmark Construction
Unseen library benchmarks are essential for evaluating retrieval-based methods in handling unseen APIs. Existing benchmarks typically involve simple API invocations or apply lexical-based evaluation metrics. To provide rigorous evaluation of complex unseen API invocations, we aim to construct execution-based multi-API benchmark that remain untrained on representative LLMs. Following prior work (Zan et al., 2024), we use Torchdata-based evaluation, which remains unexposed to powerful LLMs such as GPT-3.5 and GPT-4-0613, while allowing knowledge acquisition by newer models. This provides a valuable reference point for assessing approaches across LLMs with varying levels of API prior knowledge.
Torchdata-Manual.
We developed a new benchmark called Torchdata-Manual, comprising 100 manually crafted programming problems. Each problem involves 8-14 distinct Torchdata APIs. To ensure the diversity of the programming tasks, we randomly sampled numerous API combinations from the Torchdata documentation and selected plausible combinations to formulate the problem. Two programmers with more than five years of Python coding experience are invited to review the benchmark. More detailed construction methodology is provided in the Appendix A.4. To the best of our knowledge, Torchdata-Manual features the longest API sequences among publicly reported execution-based library-oriented benchmarks.
Torchdata-Github.
We also evaluate on an existing benchmark (Ma et al., 2024), including 50 Torchdata problems adapted from client project of Torchdata on GitHub, featuring coarse-grained user requirements that entails 3-8 API invocations. We curated the dataset by manualy supplementing external resources needed to run test cases in some problems [^1] Some external resources, such as local files to be loaded in problems, are not provided by Ma et al. (2024). and named it as Torchdata-Github.
MonkBeatEval.
To test generalizability beyond Torchdata-based evaluation, we also adapted an existing multi-library benchmark for unseen settings, with results reported in Appendix A.5.
@@ 5.1 Experimental setups
Benchmarks and base language models.
We evaluate ExploraCoder on Torchdata-Github and Torchdata-Manual benchmarks. Based on the the publicly available information on models’ training data cutoff date, we conduct our main experiments under two base models settings: (1) API-untrained model, where the API knowledge is unseen by model during training phase. We choose GPT-3.5-turbo-0125 and GPT-4-0613 as representatives. (2) API-pretrained model, where the API knowledge is pretrained in model. We represent it by GPT-4-1106-preview and two SOTA opensource code LLM: CodeQwen-1.5 and DeepseekCoder-6.7b. Due to the token budgets, we primarily experiment ExploraCoder with GPT-3.5-turbo-0125, while reporting GPT-4-0613 results where necessary to further support our conclusions.
Evaluation metrics.
We adopt Pass@k as our primary evaluation metrics. For each problem, we randomly sample $n\geq k$ code solutions from the model to execute against test cases. And pass@k is calculated as the percentage of problems solved using k candidates. To better observe nuance differences in harder problems, we additionally report Success@k (Chen et al., 2024) which relaxes the evaluation criteria by measuring whether the generated code can be executed successfully without runtime errors within limited timeout constraints.
Implementation details.
We implement ExplorCoder by setting $k_{\mathcal{D}}=4$ for task planning. For API recommendation, we set $k=20$ as initial retrieval volume, $k_{G}=15$ on Torchdata-Github following Ma et al. (2024) and $k_{G}=20$ on Torchdata-Manual. For CoAE, we set $m=5$. To generate diverse candidates, we set the $temperature=0.8$ and $top\_p=0.95$ for our CoAE and final solution generation across all baselines. More detailed experimental settings are left in Appendix A.8
$k=1$ $k=5$ $k=10$ $k=20$
API Knowledge Method Pass Success Pass Success Pass Success Pass Success
Torchdata-Github
@@
Torchdata-Manual
@@
GPT-4-1106-preview 0.16% 1.37% 0.71% 6.28% 1.62% 11.56% 2.79% 20.89%
+ naive RAG 3.19% 6.38% 12.15% 22.15% 18.30% 31.46% 24.11% 39.11%
+ ExploraCoder 14.62% 32.77% 31.19% 57.03% 37.56% 63.47% 42.20% 67.73%
@@
+ naive RAG 0.19% 0.615% 0.89% 2.92% 1.66% 5.475% 2.81% 9.53%
+ ExploraCoder 7.00% 14.80% 11.54% 22.89% 13.84% 25.40% 15.67% 27.56%
GPT-4-0613 0% 0.05% 0% 0.23% 0% 0.465 0% 0.93%
+ naive RAG 1.12% 2.94% 3.37% 8.66% 4.68% 11.98% 6.67% 16.36%
+ ExploraCoder 16.49% 24.16% 26.10% 36.89% 29.41% 40.68% 33.32% 44.32%
Table 1: Evaluation of LLMs with varying levels of prior API knowledge. We apply document-retrieval to augment the API-untrained models across two datasets, and the underperforming API-pretrained GPT4 on Torchdata-Manual.
5.2 Multi-API invocations using LLMs with varying prior API knowledge
We consider pretraining and document-retrieval as two API knowledge integration paradigms, and analyze their effectiveness in complex multi-API generation task in Table 1.
Invoking APIs using API-untrained and API-pretrained models.
By analyzing the direct generation performance of the five base models, we observe that API-pretrained models consistently outperform API-untrained models. This highlights the importance of prior API knowledge in library-oriented code generation. And the lower performance across all models on the Torchdata-Manual further underscores the challenge posed by more complex API invocations, making it a more effective benchmark for evaluation.
@@ 5.2 Multi-API invocations using LLMs with varying prior API knowledge
Through a naive RAG framework (Zhou et al., 2023), the performance of API-untrained models has been effectively improved, bridging the gap caused by the lack of prior API knowledge. We make an indirect comparison of retrieval and pretraining methods by looking into two GPT4 models (fairness dicussed in Appendix A.7). GPT-4-0613 + naive RAG outperforms GPT-4-1106-preview by an average of 6.13% pass/success rate increase on Torchdata-Github, and achieves comparable performance on more challenging Torchdata-Manual.
ExploraCoder vs naive RAG on API-untrained models.
@@
From Table 1, we can observe that ExploraCoder brings substantial improvements over naive RAG for both API-untrained models (GPT-3.5-turbo-0125 and GPT-4-0613), with an average absolute gains in pass@20 of 3.5% on Torchdata-Github and 19.8% on Torchdata-Manual. These improvements could be attributed to ExploraCoder’s potential in addressing two limitations of the naive RAG framework when handling complex API invocation subtasks:
@@
(1) Retrieval for complex requirement: In the naive RAG approach, we empirically find the retriever’s ability to recall relevant APIs for comprehensive requirements becomes a bottleneck. ExploraCoder addresses this by adopting a divide-and-conquer strategy, identifying APIs for each explicit subtask. Additionally, ExploraCoder alleviates the need for manual hyperparameter tuning by fixing retrieval counts per subtask and dynamically adjusting subtask numbers.
@@
Direct Generation 0% 0% 0% 0% 0% 0% 0% 0%
DocPrompting (2023) 0.19% 0.61% 0.89% 2.92% 1.66% 5.47% 2.81% 9.53%
CAPIR (2024) 3.01% 4.79% 6.75% 10.16% 8.21% 15.09% 9.66% 21.25%
EpiGen (2024) 2.16% 5.43% 4.40% 12.33% 5.23% 15.20% 5.86% 18.46%
ExploraCoder (Ours) 7.00% 14.8% 11.54% 22.88% 13.84% 25.4% 15.67% 27.56%
ReAct (2022) 2.00% 6.38% 2.48% 10.66% 2.95% 12.45% 3.90% 13.90%
KnowAgent (2024) 6.81% 20.54% 9.82% 22.70% 11.01% 23.29% 11.76% 23.53%
CAPIR + Self-Repair (2024) 7.47% 15.35% 8.32% 19.08% 8.64% 20.46% 8.89% 21.66%
ExploraCoder* (Ours) 11.5% 21.35% 18.32% 32.76% 20.87% 36.81% 23.51% 40.16%
Table 2: Comparing ExploraCoder with advanced retrieval-based approaches using GPT3.5 on Torchdata-Manual.
(2) Generating code with multiple unseen APIs: The complexity of coding with multiple unseen APIs lies in understanding the limited documentation and reasoning over multi-API interactions’ behavior (We provide case study in Appendix A.9). ExploraCoder mitigates this challenge by adopting a human-like exploratory programming paradigm, incrementally generating simple, reusable API invocations during CoAE, and learning extra usage knowledge from intermediate output.
ExploraCoder on API-pretrained model.
We observe the API-pretrained models underperform on Torchdata-Manual, with the most competitive GPT-4-1106-preview achieving only 0.16% in pass@1. Therefore, we use GPT-4-1106-preview on Torchdata-Manual benchmark as a proxy to further examine the effectiveness of ExploraCoder on API-pretrained models. Results in Table 1 shows ExploraCoder brings a substantial improvement for GPT-4-1106-preview, with an absolute pass@1 increase of 14.46%, and it also outperforms GPT-4-1106-preview + naive RAG by 11.43%. These results indicate that ExploraCoder is universally effective, improving models with varying levels of pretraining on relevant API knowledge.
5.3 Experience exploitation for ExploraCoder
Figure 2: Performance comparison on the Torchdata-GitHub and Torchdata-Manual datasets across two methods (ExploraCoder and ExploraCoder*). Each bar represents the mean performance GPT-3.5-turbo-0125 and GPT-4-0613 for pass/success rate, with the range lines indicating the variation between the two models.
In multi-API tasks, subsequent API invocations often depend on the outputs of earlier APIs. A failure in a dependent API could cascade into subsequent API invocations, regardless of whether their usage is correct. In ExploraCoder, unresolved subtasks could hinders the accuracy of the complete solution. An advancement of CoAE’s step-wise code generation is the API failures can be observed in early stage. This provides the opportunity to debug on the intermediate codes, which proves to outperform debugging on full code in Section 5.4.
To this end, we designed an enhanced ExploraCoder* by integrating a self-debug mechanism into CoAE. When all the candidate codes are non-executable, we exploit the failed API usage experience by debugging. Figure 2 shows ExploraCoder* significantly boosts the final solution’s quality on two models across two benchmarks, achieving an average relative increase of 55.8% in pass@1 and 71.3% in success@1. More quantitative analysis of CoAE is provided in Appendix A.3.1.
5.4 Comparing with related approaches
In this section, we further compare ExploraCoder with more advanced retrieval-based approaches. We also include Docprompting, the previously reported naive RAG framework, along with direct generation as baselines. We compare the features and computational costs in each baseline in Appendix A.1.
For SOTA multi-API-oriented methods, CAPIR and EpiGen, we set a fixed number of API recommendation in accordance with our $A_{G}$, and we directly use the subtasks generated by ExploraCoder’s planning module as the preactive plannings for EpiGen. Tables 2 and Table 3 show that ExploraCoder surpasses these methods by enriching the API knowledge from trial executions, compensating for the potential ambiguity in retrieved docs, achieving an absolute increase of 10.87% in pass@10 across the two benchmarks.
@@ 5.4 Comparing with related approaches
Direct Generation 1.70% 2.09% 5.54% 6.95% 7.28% 9.64% 8.00% 11.90%
DocPrompting (2023) 6.00% 10.57% 10.55% 24.00% 14.67% 32.50% 20.83% 40.81%
CAPIR (2024) 5.90% 10.47% 14.59% 27.08% 18.60% 37.19% 23.52% 47.43%
EpiGen (2024) 8.57% 18.95% 14.63% 35.61% 17.24% 41.67% 19.61% 47.62%
ReAct (2022) 10.19% 27.90% 10.95% 33.06% 11.90% 33.88% 13.81% 34.00%
KnowAgent (2024) 14.67% 25.81% 15.99% 30.68% 16.00% 31.90% 16.00% 33.81%
CAPIR + Self-Repair (2024) 16.47% 22.10% 21.04% 29.70% 21.75% 32.20% 22.00% 33.90%
ExploraCoder* (Ours) 19.24% 38.66% 25.41% 54.93% 27.64% 59.56% 31.62% 63.71%
Table 3: Comparing ExploraCoder with advanced retrieval-based approaches using GPT3.5 on Torchdata-Github.
To compare ExploraCoder* with other debug-enhanced methods, we first adapted a SOTA debugging framework, Self-Repair [^2] To ensure the fairness in debug iteration budget, for each problem, if ExploraCoder generates $n$ plans, enabling debugging in up to $n$ CoAE steps, we set the iteration budget for Self-Repair in that problem to $n$ accordingly., by augmenting the LLM with API knowledge retrieved by CAPIR throughout its iterative 2-stage generation. We also compare with two agentic framework, ReAct and KnowAgent, specifically designed for reactive knowledge retrieval during the solution generation. We sample the candidates from their ‘Finish‘ action, which derives final solutions based on the interleaving retrieval and debugging trajectory.
Table 2 and Table 3 shows that while these iterative/agentic methods benefit from the debugging, the overall improvement, especially in bigger k, remains limited. This could be due to the limitation of reactive planning, which is bug-driven (Appendix A.1) and lacks systematic understandings of API knowledge for diverse solution implementations. ExploraCoder*, through intermediary debugging on simpler subtasks, exhibits a significant pass@10 increase over 9.06%. Even ExploraCoder achieves a comparable performance on Torchdata-Github and surpasses them on the more complex Torchdata-Manual. This highlights our superior design in uniquely enforcing a step-wise code construction workflow and iteratively enhancing the preactive plans with exploratory knowledge.